OpenStudy (anonymous):

how do you configure 5m^-1 divided by (3m)^2 x n^-2

OpenStudy (anonymous):

\[\frac{ 5m^-1 }{ (3m)^2*n^-2 }\]

OpenStudy (anonymous):

\[\frac{ 5m^{-1} }{ 3^2*m^2*n^-2 }=\frac{ 5m^{-1} }{ 3^2*m^2*n^-2 }\]

OpenStudy (anonymous):

\[\frac{ 5m^{-1} }{ 9*m^2*n^{-2} }=\frac{ 5m^{-1-2} }{ 9*n^{-2} }\] \[\frac{ 5m^{-3} }{ 9*n^{-2} }\]